OpenStudy (anonymous):

Algebra 2 √(x+14) = x+8 (find x ) (the square root of x plus 14 is equal to x plus 8)

OpenStudy (anonymous):

Square both sides to eliminate the square root x+14=(x+8)^2

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Now expand the right side

OpenStudy (anonymous):

what does that mean??? can you show me by drawing??

OpenStudy (anonymous):

(x+8)^2 is (x+8)(x+8) can you foil that?

OpenStudy (anonymous):

yes.... x^2 + 16x + 64

OpenStudy (anonymous):

Right so now we have x+14=x^2+16x+64 move the x and 14 to the right side

OpenStudy (anonymous):

so.... x^2 + 15x + 50

OpenStudy (anonymous):

Yes, can you factor that?

OpenStudy (anonymous):



OpenStudy (anonymous):

( x + 5)(x + 10)

OpenStudy (anonymous):

So x=? and x=?

OpenStudy (anonymous):

x = -5 x = - 10

OpenStudy (anonymous):

You'll need to check solutions since theres a square root involved

OpenStudy (anonymous):

So plug each one of those in for x, do they both work?

OpenStudy (anonymous):

 no